# Data Compression for Fast Online Stochastic Optimization

### v1

@@ 3.3 Clustered distribution analysis
Since the performance bounds will be affected by the sequence of clustered empirical distributions $\hat{{\mathbf{P}}}^{K}_{t}$, we first derive some results pertaining these distributions. In particular, we formalize its distances, under different metrics, to the empirical distribution $\hat{{\mathbf{P}}}_{t}$, as well as its convergence to a limiting distribution ${\mathbf{P}}^{K}_{\star}$. We define the following values.
@@
For any clustered configuration of the dataset $\mathcal{D}_{t}$, with clusters $C^{k}_{t}$, cluster centroids $\bar{u}^{k}_{t}$, and clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we define the Wasserstein cost $d^{K}_{t,p}$ and clustering value $D^{K}_{t,p}$ as
$d^{K}_{t,p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t}),\quad D^{K}_{t,p}=\left(\frac{1}{n_{t}}\sum_{k=1}^{K_{t}}\sum_{\hat{u}\in C^{k}_{t}}\left\|\hat{u}-\bar{u}^{k}_{t}\right\|_{2}^{p}\right)^{1/p}.$
@@
where $\mathbf{1}\{\cdot\}$ is the 00-$1$ indicator function. Note that $\Phi^{K}_{t}=0$ when $J=1$, i.e., when $f$ is a concave function.
By construction, the above values all quantify the distance between the clustered and the non-clustered empirical distributions, under different metrics, and are nonnegative. In addition, we establish a hierarchy between these distances and provide a few bounds, making use of the following definition.
@@
[…]
@@
By definition, the Wasserstein distance is computed with the optimal coupling (joint distribution) between the two reference distributions, while $D^{K}_{t,p}$ is calculated from one particular coupling. Therefore, $d^{K}_{t,p}\leq D^{K}_{t,p}$. The lower bound on $d^{K}_{t,p}$ follows from Proposition 3.1, and the equalities follow from the fact that $\hat{{\mathbf{P}}}^{K}_{t}=\hat{{\mathbf{P}}}_{t}$ when $K\geq n_{t}$. The final inequalities follow from the boundedness of the support. ∎
Furthermore, we can derive asymptotic bounds, for $t\to\infty$. This relies on Assumption 3.2, that for $t\geq\tau$, where $\tau<\infty$ is some finite time, the cluster supports $S_{t}^{k}$ are fixed. We make use of the following lemma.
Lemma 3.1 (Clustering convergence).
Under Assumption 3.2 on the cluster supports, the corresponding sequence of clustered empirical distributions $\hat{{\mathbf{P}}}^{K}_{t}$ converges almost surely with respect to ${\mathbf{P}}^{\infty}$ to a distribution ${\mathbf{P}}^{K}_{\star}$. Furthermore, the distributions achieve almost sure convergence with respect to the Wasserstein metric, that is,
@@
Without loss of generality, let $K_{\tau}=K$. By Assumption 3.2, the cluster supports are fixed. By the Strong Law of Large Numbers (SLLN) [34], for each cluster $k\leq K$,
@@
This shows that the cluster weights converge almost surely to ${\mathbf{P}}(S^{k}_{\tau})=\theta_{\star}^{k}$, the true weights on each cluster support.
Since, for all clusters $k$, ${\mathbf{P}}(S^{k}_{\tau})>0$, we note that $n^{k}_{t}\to\infty$ when $t\to\infty$. Furthermore, since ${\mathbf{P}}$ has finite $p$-th moments, the conditional distributions of ${\mathbf{P}}$ on these supports $S^{k}_{\tau}$ must also have finite $p$-th moments. We denote these conditional distributions ${\mathbf{P}}^{k}$, and their means $\bar{u}^{k}_{\star}$. We then observe, by the SLLN,
@@
i.e., the cluster centroids converge almost surely to the true means of the conditional distributions ${\mathbf{P}}^{k}$.
By Assumption 3.2, the cluster supports are pairwise disjoint, and cover the entire support $S$. It follows that $\sum_{k=1}^{K}{\mathbf{P}}(S^{k}_{\tau})=1$. This implies
@@
where ${\mathbf{P}}^{K}_{\star}$ is a discretized distribution of ${\mathbf{P}}$ onto $K$ atoms. Since ${\mathbf{P}}^{K}_{\star}$ has finite $p$-th moments by construction, this implies almost sure convergence with respect to the Wasserstein-$p$ metric [55, Theorem 6.9]. ∎
@@
Under the Lemma 3.1, we have
$\mathbf{P}^{\infty}\left\{d^{K}_{\star,p}({\mathbf{P}})\leq\lim_{t\rightarrow\infty}d^{K}_{t,p}\leq W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}=1.$
Furthermore, as $K\to\infty$, $\lim_{K\rightarrow\infty}W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})=0$ for all $p$.
@@
By the triangle inequality, we note that for all $t$,
$d^{K}_{t,p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t})\leq W_{p}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})+W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+W_{p}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t}),$
that is, the distance between the two empirical distributions is bounded by the sum of the distances between the nominal empirical distribution and ${\mathbf{P}}$, between ${\mathbf{P}}$ and the converged clustered distribution, and between converged clustered distribution and the clustered empirical distribution. Taking the limit as $t\to\infty$, we have
$\lim_{t\to\infty}d^{K}_{t,p}\leq\lim_{t\to\infty}W_{p}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})+W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+\lim_{t\to\infty}W_{p}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t}).$
Under the given assumptions, Theorem 3.1 implies
$\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}W_{p}(\mathbf{P},\hat{{\mathbf{P}}}_{t})=0\right\}=1.$ (11)
Together with Lemma 3.1, we obtain $\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}d^{K}_{t,p}\leq W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}=1.$ For the lower bound, recall that by Proposition 3.1, $d^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})\leq d^{K}_{t,p}.$ The result then holds by the asymptotic consistency of finite quantizations, given in [24, Corollary 4.24], which holds due to (11). Specifically, we note that $\mathbf{P}^{\infty}\left\{\lim_{t\rightarrow\infty}d^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})=d^{K}_{\star,p}({\mathbf{P}})\right\}=1.$ Lastly, the asymptotic result for $K\to\infty$ holds from [24, Lemma 6.1]. ∎
The asymptotic limits of $D^{K}_{t,p}$ and $\Phi^{K}_{t}$ follow similarly.
@@
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}D^{K}_{t,p}=D^{K}_{\star,p}=\left(\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\|u-\bar{u}^{k}_{\star}\|^{p}_{2}\right]\right)^{1/p}\right\}.$
Furthermore, with accumulation points $\hat{z}^{jk}_{\star}=\lim_{t\to\infty}z_{t}^{jk}$, we have
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\Phi^{K}_{t}=\Phi^{K}_{\star}=\mathbf{1}\{J\geq 2\}\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\max_{j\leq J}(z^{jk}_{\star})^{T}(u-\bar{u}^{k}_{\star})\right]\right\}.$
@@
This follows from the fixed cluster supports and the almost-sure convergence of both empirical distributions. ∎
From Lemma 3.1, we established that the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$ converges to the distribution ${\mathbf{P}}^{K}_{\star}$. We can also prove that its convergence rate follows the convergence rate of the non-clustered empirical distribution.
@@
Suppose Assumption 3.2 holds, and we have a sequence $\varepsilon_{t}>0$, computed using Theorem 3.1, such that $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$ and the corresponding confidence sequence $\beta_{t}\in(0,1)$ satisfies $\sum_{t=1}^{\infty}\beta_{t}<\infty$. Then, for $t\geq\tau$, with rate $O(\varepsilon_{t})$, the sequence of centroids $\bar{u}^{k}_{t}$ converges to the true centroids $\bar{u}^{k}_{\star}$, and the distance $W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})$ converges to 0.
@@
By assumption, all cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{\tau})>0$. Since ${\mathbf{P}}$ is light-tailed, we also know that for some $r>1$, there is some finite $R<\infty$ defined by Assumption 3.1. It follows that the conditional distributions of ${\mathbf{P}}$ on the supports $S^{k}_{\tau}$, denoted ${\mathbf{P}}^{k}$, are also light-tailed, with the same $r$ and a constant $R^{k}\leq O(R)$, i.e., $R^{k}=\int_{S_{\tau}^{k}}\exp(\|u\|^{r}){\mathbf{P}}(u|u\in S^{k}_{\tau})(du)<\infty.$ By Theorem 3.1, we have
${\mathbf{P}}^{n_{t}}(W_{p}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{t,k})\leq O(\varepsilon_{t}))\geq 1-\beta_{t},$
@@
This implies that the centroids $\bar{u}^{k}_{t}$ converge to the true means $\bar{u}^{k}_{\star}$ with rate $O(\varepsilon_{t})$, with respect to the infinity norm, i.e., $\|\bar{u}^{k}_{t}-\bar{u}^{k}_{\star}\|_{\infty}\leq O(\varepsilon_{t}).$ Then, by the ordering of norms and the definition of the Wasserstein distance on discrete distributions,
@@ 3.4 Performance guarantees
Assumption 3.3 (Ambiguity set construction).
Let the following hold.
@@
    For all $t$, $\varepsilon_{t}$ is computed using Theorem 3.1, with a sequence $\beta_{t}$ such that $\sum_{t=0}^{\infty}\beta_{t}<\infty$ and $\lim_{t\to\infty}\varepsilon_{t}=0$.
@@
    The adaptive ambiguity set $\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{p}(\hat{\mathbf{P}}^{K}_{t})$, with $\varepsilon^{K}_{t}=\varepsilon_{t}$.
In addition, either the Lipschitz or smoothness condition must hold.
Assumption 3.4 (Lipschitzness and smoothness).
Let at least one hold.
  1. 1.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the Lipschitz condition given in Definition 2.1, with constants $M_{j}$.
  2. 2.
    For all $x\in\mathcal{X}$, the constituent functions $f_{j}$ of $f$ satisfy the smoothness condition given in Definition 2.2, with constants $L_{j}$.
We present the main results after the following lemma. Proofs are in Appendix A.1.
Lemma 3.2 (Effect of the support $S$).
Let $\Delta_{t}=\tilde{H}_{t}-H_{t}$, where $\tilde{H}_{t}$ is the value of the nominal DRO problem where the support $S={\mbox{\bf R}}^{d}$, i.e., the ambiguity set is defined $\tilde{\mathbf{B}}_{\varepsilon_{t-1}}^{p}(\hat{{\mathbf{P}}}_{t-1})=\left\{\mathbf{Q}\in\mathcal{W}({{\mbox{\bf R}}^{d}})\mid W_{p}(\hat{{\mathbf{P}}}_{t-1},\mathbf{Q})\leq\varepsilon_{t-1}\right\},$ and the solution is fixed at $x_{t}$. If the sequence of $\varepsilon_{t}$ decreases following clause 1 of Assumption 3.3, we have ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\Delta_{t}=0\right\}=1$.
This lemma formalizes the discrepancy between nominal DRO objectives with and without explicit support constraints. The discrepancy appears in the following bound, but we note that it converges to 0 as $\varepsilon_{t}\rightarrow 0$, and is often negligible.
Theorem 3.5 (Finite sample guarantee).

### v2

@@ 3.2 Clustered distribution analysis
Since the performance bounds will be affected by the sequence of clustered empirical distributions $\hat{{\mathbf{P}}}^{K}_{t}$, we first derive some results pertaining these distributions. In particular, we formalize its distances, under different metrics, to the empirical distribution $\hat{{\mathbf{P}}}_{t}$, as well as prove its convergence to a limiting distribution ${\mathbf{P}}^{K}_{\star}$. We define the following values. Note that the orders and norms below are independent of the order and norm of the ambiguity set (8).
@@
For any clustered configuration of the dataset $\mathcal{D}_{t}$, with clusters $C^{k}_{t}$, cluster means $\bar{u}^{k}_{t}$, and clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$, we define the Wasserstein cost $W^{K}_{p}$ and clustering value $D^{K}_{t,p}$ as follows.
$W^{K}_{t,p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t}),\quad D^{K}_{t,p}=\left(\frac{1}{n_{t}}\sum_{k=1}^{K_{t}}\sum_{\hat{u}\in C^{k}_{t}}\left\|\hat{u}-\bar{u}^{k}_{t}\right\|_{2}^{p}\right)^{1/{p}}.$
@@
where $\mathbf{1}\{\cdot\}$ is the 0-11 indicator function. Note that $\Phi^{K}_{t}=0$ when $J=1$, i.e., when $f$ is a concave function.
By construction, the above values all quantify the distance between the clustered and the non-clustered empirical distributions; the first two are distance metrics, while the third is a nonnegative cost function. We establish a hierarchy between these distances and provide a few bounds, making use of the following definition.
@@
For all $K$, $t$, and $p$, with $\|\cdot\|$ the $\ell_{2}$-norm, $W^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})\leq W^{K}_{t,p}\leq D^{K}_{t,p}.$ Under Assumption 3.1, if $K\geq n_{t}$, $D^{K}_{t,p}=W^{K}_{t,p}=\Phi^{K}_{t}=0.$ In addition, if $S$ is bounded, with radius $\rho=(1/2)\mathrm{diam}_{\infty}(S)<\infty$, we have that for all $K$ and $t$, and any $\|\cdot\|$,
@@
The Wasserstein distance is computed with the optimal (minimizing) coupling (joint distribution) between the two reference distributions, while $D^{K}_{t,p}$ is calculated from one particular coupling. Therefore, $W^{K}_{t,p}\leq D^{K}_{t,p}$. The lower bound on $W^{K}_{t,p}$ follows by Definition 3.3, and the equalities follow from the fact that $\hat{{\mathbf{P}}}^{K}_{t}=\hat{{\mathbf{P}}}_{t}$ when $K\geq n_{t}$. The final inequalities follow from the boundedness of the support. ∎
Furthermore, we can derive asymptotic bounds, for $t\to\infty$. This relies on Assumption 3.1, that for $t\geq\tau$, where $\tau<\infty$ is some finite time, we fix the cluster supports $S_{t}^{k}$. We make use of the following lemma.
Lemma 3.1 (Clustering consistency).
[…]
@@
W.l.o.g., let $K_{\tau}=K$. By Assumption 3.1 and the Strong Law of Large Numbers (SLLN) [31], for each cluster $k\leq K$,
@@
This shows that the cluster weights converge almost surely to ${\mathbf{P}}(S^{k}_{\tau})=\theta_{\star}^{k}$, the true weights on each cluster support. Next, as ${\mathbf{P}}(S^{k}_{\tau})>0$ for all clusters $k$, we note that $n^{k}_{t}\to\infty$ when $t\to\infty$. Furthermore, since $\mathbf{P}$} has finite $p$-th moments, the conditional distributions of $\mathbf{P}$} on these supports $S^{k}_{\tau}$ must also have finite $p$-th moments. We denote these conditional distributions ${\mathbf{P}}^{k}$, and their means $\bar{u}^{k}_{\star}$. Then, by the SLLN,
@@
i.e., the cluster means converge almost surely to the true (finite) means of the conditional distributions ${\mathbf{P}}^{k}$. Since both the probabilities and the means converge almost surely to finite limits, it follows that
@@
where ${\mathbf{P}}^{K}_{\star}$ is constructed to be a discretized distribution of $\mathbf{P}$} onto $K$ atoms. Since ${\mathbf{P}}^{K}_{\star}$ has finite $p$-th moments by construction, this implies almost sure convergence with respect to the Wasserstein-$p$ metric [53, Theorem 6.9]. ∎
@@
Under Lemma 3.1, we have
$\mathbf{P}^{\infty}\left\{W^{K}_{\star,p}({\mathbf{P}})\leq\lim_{t\rightarrow\infty}W^{K}_{t,p}=W^{K}_{\star,p}=W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\right\}=1.$
Furthermore, $\lim_{K\rightarrow\infty}W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})=0$ for all $p$.
@@
By the continuity of the Wasserstein-$p$ metric, and by the almost-sure convergence of the empirical distributions to the respective true distributions (Theorem 2.1 and Lemma 3.1), we obtain $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W^{K}_{t,p}=W_{p}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})\}=1$ [52, Theorem 7.12]. For the lower bound, recall that by Definition 3.3, $W^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})\leq W^{K}_{t,p}.$ The result then holds by the asymptotic consistency of finite quantizations, given in [20, Corollary 4.24]. Specifically, we note that $\mathbf{P}^{\infty}\{\lim_{t\rightarrow\infty}W^{K}_{\star,p}(\hat{{\mathbf{P}}}_{t})=W^{K}_{\star,p}({\mathbf{P}})\}=1.$ Lastly, the asymptotic result for $K\to\infty$ holds from [20, Lemma 6.1]. ∎
The asymptotic limit of $D^{K}_{t,p}$ and the limit superior of $\Phi^{K}_{t}$ follows similarly.
@@
${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}D^{K}_{t,p}=D^{K}_{\star,p}=\left(\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\|u-\bar{u}^{k}_{\star}\|^{p}_{2}\right]\right)^{1/{p}}\right\}=1.$
Furthermore, if there exist accumulation points of $z_{t}^{jk}$ for all $j\leq J,k\leq K$, then
${\mathbf{P}}^{\infty}\left\{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\limsup_{t\to\infty}}~\Phi^{K}_{t}=\Phi^{K}_{\star}=\mathbf{1}\{J\geq 2\}\mathbf{E}_{{\mathbf{P}}}\left[\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\max_{j\leq J}(-z^{jk}_{\star})^{T}(u-\bar{u}^{k}_{\star})\right]\right\}=1,$
where $\hat{z}^{jk}_{\star}$ are the accumulation points that achieve the limit superior.
@@
By the consistency of our clustering procedure (Lemma 3.1), the 0-measure of the cluster boundaries (Assumption 3.1), the continuity of linear functions and $\|\cdot\|^{p}$, and the existence of the accumulation points, the limit (or limit superior, for the second result) of the functions within the summations of Definition 3.4 converge almost surely to the functions within the expectations above, for $\mathbf{P}$}-almost every $u$. That is, defining $\omega_{t}(u)=(\sum_{k=1}^{K_{t}}\mathbf{1}\{u\in C^{k}_{t}\}\|u-\bar{u}^{k}_{t}\|_{2}^{p})^{1/{p}}$ for the first result, by the given assumptions, we have that $\lim_{t\to\infty}\omega_{t}(u)=\sum_{k=1}^{K}\mathbf{1}\{u\in S^{k}_{\tau}\}\|u-\bar{u}^{k}_{\star}\|^{p}_{2}=\omega_{\star}(u)$, for $\mathbf{P}$}-almost every $u$. Next, to conclude that $\lim_{t\to\infty}\mathbf{E}_{\hat{{\mathbf{P}}}_{t}}[\omega_{t}(u)]=\mathbf{E}_{{\mathbf{P}}}[\omega_{\star}(u)]$, we require the function $\omega_{t}(u)$ to have an integrable dominating function. W.l.o.g., let $u$ belong in the cluster $k$. By the triangle inequality, $|\omega_{t}(u)|=\|u-\bar{u}^{k}_{t}\|^{p}_{2}\leq(\|u\|_{2}+\|\bar{u}\|^{k}_{t})^{p}\leq\bar{C}(1+\|u\|_{2}^{p})$ for some constant $\bar{C}$. By the assumption of finite $p$-th moments, this dominating function is integrable. The result then follows from [52, Theorem 7.12(iv)], which establishes a generalized convergence of expectation. Since we have shown almost sure pointwise convergence in the first step, the cited theorem holds even without the assumption on continuity.
From Lemma 3.1, we established that the clustered empirical distribution $\hat{{\mathbf{P}}}^{K}_{t}$ converges almost surely to the distribution ${\mathbf{P}}^{K}_{\star}$. We can also prove that its convergence rate follows the convergence rate of the non-clustered empirical distribution.
@@
Suppose Assumption 3.1 holds, and we have a sequence $\varepsilon_{t}$, computed using Theorem 2.1 for a order $p$, such that $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$ and the confidence sequence $\beta_{t}\in(0,1)$ satisfies $\sum_{t=1}^{\infty}{\beta_{t}<\infty}$. Then, for $t\geq\tau$, the distance $W_{p}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})$ converges to 0 with rate $O(\varepsilon_{t})$.
@@
By assumption, all cluster supports have nonzero measure, i.e., ${\mathbf{P}}(S^{k}_{\tau})>0$. Since $\mathbf{P}$} is light-tailed, we also know that for some $\alpha>1$, there is some finite $R<\infty$ defined by Assumption 2.3. It follows that the conditional distributions of $\mathbf{P}$} on the supports $S^{k}_{\tau}$, denoted ${\mathbf{P}}^{k}$, are also light-tailed, with the same $\alpha$ and a constant $R^{k}\leq O(R)$, i.e., $R^{k}=\int_{S_{\tau}^{k}}\exp(\|u\|^{\alpha}){\mathbf{P}}(u|u\in S^{k}_{\tau})(du)<\infty.$ By Theorem 2.1,
${\mathbf{P}}^{t}(W_{p}({\mathbf{P}}^{k},\hat{{\mathbf{P}}}^{t,k})\leq O(\varepsilon_{t}))\geq 1-\beta_{t},$
@@
This implies that the means $\bar{u}^{k}_{t}$ converge to the true means $\bar{u}^{k}_{\star}$ with rate $O(\varepsilon_{t})$, with respect to the infinity norm, i.e., $\|\bar{u}^{k}_{t}-\bar{u}^{k}_{\star}\|_{\infty}\leq O(\varepsilon_{t}).$ Then, by the equivalence of norms on ${\bf R}^{d}$ and the definition of the Wasserstein distance on discrete distributions,
@@ 3.3 Performance guarantees
Assumption 3.2.
Fix an order $r$ and let the following hold.
@@
    For the chosen order $r$ and all $t$, $\varepsilon_{t},\beta_{t}$, and $\rho_{t}$ are constructed according to either Theorem 2.1 or Theorem 2.2, under their respective assumptions.
@@
    The adaptive ambiguity set $\mathcal{P}^{K}_{t}=\mathbf{B}_{\varepsilon^{K}_{t}}^{r}(\hat{\mathbf{P}}^{K}_{t})$, with $\varepsilon^{K}_{t}=\varepsilon_{t}$ and the Wasserstein distance defined with an arbitrary norm $\|\cdot\|$.
3.3.1 Concave case
We first present a simplified set of results for a concave objective function $f$, with unbounded support $S={\bf R}^{d}$. The proofs can be derived from the general case, which will be presented subsequently.
Theorem 3.4 (Concave finite sample guarantee).